Please answer the following physics questions:

1.) Consider a simple pendulum (no friction). Does the angular acceleration of the mass remain constant during its motion? If not, then how does the angular acceleration change? In which part of the motion, the angular acceleration of the pendulum would be equal to zero?

2.) A pendulum has th period equal to 1,000 s on the Earth. What would be the period of the same pendulum when measured on the surface of the moon? Would the period be the same as it is on the Earth?
Answer #1

No angualr accleration is not constant it is sinusoidal with amplitude g/L

it is zero whenver sin (wt) = 0, at which time

a' = sqrt g/L, that is exactly when a= 0

(2)

T _E/T_M = sqrt g_M/g_E

1000 s/T_M = sqrt 3.711/9.81

T_M = 1625.88 s